1. You deposit $100,000 cash in a brokerage account and short sell $200,000 of stocks on margin. Later, the value of the stocks held short rises to $250,000. What is your account margin in percent?

1. You deposit $100,000 cash in a brokerage account and purchase $200,000 of stocks on margin by borrowing $100,000 from your broker. Later, the value of your stock holdings falls to $150,000. What is your account margin in percent?
Answer #1

1). Given, Broker's fund = $100000

Own fund = $100000

if stock price rise to $250000

as the stock is short sell, it will give a loss of 250000-200000 = $50000

So new own fund = $100000-50000 = $50000

So, Account margin in percent = 50000/150000 = 33.33%

2).

1). Given, Broker's fund = $100000

Own fund = $100000

if stock price falls to $150000

as the stock is purchased, it will give a loss of 200000-150000 = $50000

So new own fund = $100000-50000 = $50000

So, Account margin in percent = 50000/150000 = 33.33%